A 4-person grievance committee is to be selected out of 2
departments, A and B, with 17 and 29 people. In how many
ways can the following committees be selected.
A) 3 from A and 1 from B
B) 2 from A and 2 from B
C) All from A
D) 4 people regardless of department
E) at least 3 from department A.

Answer by Edwin McCravy(20055)

A 4-person grievance committee is to be selected out of 2
departments, A and B, with 17 and 29 people. In how many
ways can the following committees be selected.
A) 3 from A and 1 from B
`
(17C3)(29C1) = (680)(29) = 19720
`
`
B) 2 from A and 2 from B
`
`
(17C2)(29C2) = (136)(406) = 55216
`
`
C) All from A
`
`
(17C4)(29C0) = (2380)(1) = 2380
`
`
D) 4 people regardless of department
`
`
(17+29)C4 = 46C4 = 163185

E) at least 3 from department A.

(17C3)(29C1) + (17C4)(29C0) =
(680)(29) + (2380)(1) = 19720 + 2380 = 22100